### v2

@@ Appendix C Missing Proofs of Section 4
Appendix D MMS with copies for $k$-demand valuations
It is natural to consider other classes of valuation functions, not only arbitrary monotone and additive. One of the simplest valuations is that of 1-demand, every agent only wants one good, and every good has a value. Trivially, one round of round-robin will result in an MMS allocation. So, what about $k$-demand? The valuation of a set is the sum of the $k$ most valuable goods in this set.
It is natural to ask if the algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ is relevant, unfortunately, this does not work. Consider the following input: All $n$ agents with identical valuation functions $v_{i}=v$, $v$ is 2-demand: $v(g)\geq 0$ for all goods, and $v(S)=\max_{g,h\in S}v(g)+v(h)$ for $|S|\geq 2$.
We now construct an example on which algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ fails, we define goods and valuations as follows: there are $n-1$ goods $g_{1},\ldots,g_{n-1}$ for which $v(g_{j})=(1-\epsilon)/2$, $n-1$ goods $h_{1},\ldots,h_{n-1}$ for which $v(h_{j})=(1+\epsilon)/2$, and two additional goods $x_{1}$, $x_{2}$ for which $v(x_{j})=1/2$. One possible MMS allocation is $\{g_{1},h_{1}\}\ldots\{g_{n-1},h_{n-1}\},\{x_{1},x_{2}\}$, and the MMS value is one.
Now consider [arbitrarily] adding goods in the order $g_{1},\ldots,g_{n-1}$ to a bag $B$, at this point in time $v(B)<1$. Next, add $x_{1}$ and $x_{2}$ to $B$, all agents have value 1 for the bag, if the bag is assigned to one of the agents not enough goods will remain. Duplicating the good $x_{2}$ and all of the $h_{j}$ will not suffice to get an MMS allocation for the remaining $n-1$ agents.
This suggests a variant of algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ for $k$-demand valuations. Upon processing a new good $g$, and adding it to $B$, do not assign the entire bag $B$ to an agent with $v(B)\geq 1$. Instead, choose some set $S\subseteq B$, $|S|=k$ with $v(S)\geq 1$. Some such set $S$ must exist (as $v(B)\geq 1$ and $v$ is $k$-demand). Moreover, the last good added to the bag $g$ must be in $S$, otherwise the bag $B^{\prime}=B\setminus\{g\}\subset B$, considered just prior to the arrival of $g$, would also have had value $\geq 1$. Give one of the agents with $v(B)\geq 1$ some $k$-subset of $B$, $S\subset B$, for which $v(S)\geq 1$. Return all goods in $B\setminus S$ to the pool of goods yet to be processed, and assign $B=\{g\}$, the last good added to that triggered this allocation, this creates a copy of $g$. Also, such duplications can be done at most $n-1$ times and no good is duplicated more than once.
To Summarize:
Observation 4.
For $k$-demand valuations $n-1$ distinct copies suffice to achieve an MMS allocation.
Appendix E Example for Algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$
@@
Consider the case when we add items in order $g_{4},g_{7},g_{9},g_{1},g_{2},g_{10},g_{12},g_{7},g_{5},g_{3},g_{8}\ldots$
@@
  • For $B=\{g_{4},g_{7},g_{9}\}$, $v_{1}(B)=1.2$, $v_{2}(B)=0.8$, $v_{3}(B)=1.1$, and $v_{4}(B)=1.1$, arbitrarily choose agent 1 and assign $B$ to agent 1, set $B$ to be a copy of the last good added: $B=\{g_{9}\}$.
  • For $B=\{g_{9},g_{1},g_{2}\}$, we have $v_{2}(B)=1.1$, $v_{3}(B)=0.8$, $v_{4}(B)=1$, arbitrarily choose agent 2 and assign $B$ to agent 2, set $B=\{g_{2}\}$.
  • For $B=\{g_{2},g_{10},g_{12},g_{7}\}$, we have $v_{3}(B)=1.2$, $v_{4}(B)=1$ and assign $B$ to agent 4, set $B=g_{7}$.
  • For $B=\{g_{7},g_{5},g_{3},g_{8}\}$ we have that $v_{3}(B)=1.5$ and assign $B$ to agent 3.
@@
Appendix F Picking Sequence Fails for Allocation with Copies
@@
Consider the ordered counterpart (see Definition 4) of this instance and the allocation $A^{\prime}$ below, in which every good is copied once.
@@
One can generalize the picking sequence algorithm ([16], see also Proposition 2), to a setting with copies in two ways. Assume we have a virtual allocation $A^{\prime}$, where each good has at most $k$ extra copies.
One way is to run $k+1$ instances of the picking sequence algorithm, where in each run each good in $M$ is allocated at most once. The first agent to pick is an agent which holds (one of the instances of) the virtual good $1$, the second holds $2$ and so on. Another way, is to allow all agents which hold good $1$ to pick first, then all agents which hold good $2$, and so on. In this interpretation agent $i$ picks the most-valued good $g$ such that $g\notin A_{i}$, and $g$ was allocated no more than $k+1$ times (so each original good is copied at most $k$ times as well). Below we show that neither of these interpretations work if we break the ties arbitrarily. In fact, the first one fails for any tie-breaking rule, see Example 2.
The failure of the two generalizations follows from the fact that an agent is not allowed to have two instances of the same good. If valuations were such that any $q$ copies of any good $g$ would yield a value of $q\cdot v(g)$, the same proof as in [16] would carry over to this setting. We show examples in which the two generalizations fail when at most one extra copy is made (i.e., for $k=1$). For the first generalization, the example below shows that breaking ties arbitrarily may fail.
Example 1.

### v3

@@ Appendix B Missing Proofs of Section 3
B.1 Example for Algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$
@@
Consider the case when we add items in order $g_{4},g_{7},g_{9},g_{1},g_{2},g_{10},g_{12},g_{6},g_{5},g_{3},g_{8},g_{11}$.
@@
  • For $B=\{g_{4},g_{7},g_{9}\}$, $v_{1}(B)=1.2$, $v_{2}(B)=0.6$, $v_{3}(B)=1.1$, and $v_{4}(B)=1.1$, arbitrarily choose agent 1 and assign $B$ to agent 1, set $B$ to be a copy of the last good added: $B=\{g_{9}\}$.
  • For $B=\{g_{9},g_{1},g_{2}\}$, we have $v_{2}(B)=1.1$, $v_{3}(B)=1.4$, $v_{4}(B)=1$, arbitrarily choose agent 2 and assign $B$ to agent 2, set $B=\{g_{2}\}$.
  • For $B=\{g_{2},g_{10},g_{12},g_{6}\}$, we have $v_{3}(B)=0.6$, $v_{4}(B)=1.6$ and assign $B$ to agent 4, set $B=g_{6}$.
  • For $B=\{g_{6},g_{5},g_{3},g_{8},g_{11}\}$ we have that $v_{3}(B)=2$ and assign $B$ to agent 3.
@@
Appendix C Missing Proofs of Section 4
See 4.12
Proof.
See 4.9
Proof.
Let $P=(P_{1},P_{2},\ldots,P_{n})$ be an MMS partition of agent $i$ for the original instance. By the pigeonhole principle, there exists $j\in[n]$ such that $|P_{j}\cap[kn+1]|>k$. Without loss of generality, let us assume $|P_{n}\cap[kn+1]|>k$ and $g_{1},g_{2},\ldots g_{k+1}$ are $k+1$ distinct goods in $P_{n}\cap[kn+1]$ such that $g_{1}\leq g_{2}\leq\ldots\leq g_{k+1}$. For $j\in[k+1]$, let $kn-k+j\in P_{a_{j}}$. Now for all $j\in[k+1]$, swap the goods knk+jkn-k+j and $g_{j}$; i.e., iteratively do the following:
Let $P^{\prime}=(P^{\prime}_{1},P^{\prime}_{2},\ldots,P^{\prime}_{n})$ be the final partition. Since $g_{1},\ldots,g_{k+1}$ are $k+1$ goods in $[kn+1]$ in increasing order of index, $g_{j}\leq kn-k+j$. Thus, $v_{i}(g_{j})\geq v_{i}(kn-k+j)$ and after each of these swaps, the value of $P_{a_{j}}$ cannot decrease. Therefore, for all $j\in[n-1]$, $v_{i}(P^{\prime}_{j})\geq v_{i}(P_{j})\geq\mu^{n}_{i}(M)$ and $(P^{\prime}_{1},\ldots,P^{\prime}_{n-1})$ is a partition of a subset of $M\setminus K$ into n1n-1 bundles of value at least $\mu^{n}_{i}(M)$ for $i$. The lemma follows. ∎
See 4.13
Proof.
To prove the lemma, it suffices to find a partition of (a subset of) $M\setminus\{g_{1},g_{2}\}$ into n1n-1 bundles, such that the value of agent $i$ for each bundle is at least $\mu^{n}_{i}(M)$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $\{g_{1},g_{2}\}\subseteq P_{1}\cup P_{2}$. If $\{g_{1},g_{2}\}\subseteq P_{1}$, then $(P_{2},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, let $g_{i}\in P_{i}$ for $i\in[2]$. We have
See 4.14
Proof.
To prove the lemma, it suffices to partition (a subset of) $M\setminus\{g_{1},g_{2},g_{3}\}$ into n2n-2 bundles, each of value at least $\mu^{n}_{i}(M)$ to $i$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $\{g_{1},g_{2},g_{3}\}\subseteq P_{1}\cup\ldots\cup P_{j}$ for $j\leq 3$. If $\{g_{1},g_{2},g_{3}\}\subseteq P_{1}\cup P_{2}$, then $(P_{3},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, let $g_{i}\in P_{i}$ for $i\in[3]$. We have
See 4.15
Proof.
For all agents $i$ who is assigned a bundle $B$ in $S^{\alpha}$, we have $v_{i}(B)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. First consider the case that $S^{\alpha}$ removes two agents $i\neq j$. Since the reduction rule $R^{\alpha}_{1}$ is not applicable, for all the remaining agents $a$, we have $v_{a}(\{n,n+1\})<\alpha\cdot\mu^{n}_{a}(M)\leq\mu^{n}_{a}(M)$. Hence, by lemma 4.14, $\mu^{n-2}_{a}(M\setminus\{1,n,n+1\})\geq\mu^{n}_{a}(M)$.
Now consider the case that $S^{\alpha}$ removes only one agent $i$. Let $a\neq i$ be a remaining agent after allocating $\{1,n+1\}$ to $i$. We have $v_{a}(\{1,n+1\})<\alpha\cdot\mu^{n}_{i}(M)\leq\mu^{n}_{i}(M)$, otherwise, $S^{\alpha}$ would allocate $\{1,n\}$ to $a$ (or some agent other than $i$). By Lemma 4.13, $\mu^{n-1}_{a}(M\setminus\{1,n+1\})\geq\mu^{n}_{a}(M)$. ∎
See 4.3
Proof.
We prove for ordered instances $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, $\mathsf{BagFill\mbox{-}RoundRobin}(\mathtt{T-reduce(\mathcal{I},\alpha)},\alpha)$ returns a simple $\alpha$-MMS allocation with at most $\lfloor n^{*}/3\rfloor$ many distinct copies. Then, from Proposition 4.1, the theorem follows.
Let $n$ be the number of agents after applying $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Since $T^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{T-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. Without loss of generality, let the MMS value of the remaining agents be $1$ at this point of the algorithm. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$ fraction of their MMS. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. Since $g\leq n^{\prime}$, $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 4.30, $v_{i}(B_{j}\cup\{g\})\geq\alpha$. Therefore, the final allocation is $\alpha$-MMS.
We now give an upper bound on the total number of copies and prove that the generated MMS allocation is simple, namely that there are $t\leq\lfloor n^{*}/3\rfloor$ distinct copies, and that the $2t$ instances, composed of these copies and their corresponding original goods, are allocated to $2t$ distinct agents. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Only when applying $T^{\alpha}$, new duplications are introduced and one duplication per three agents that are removed. So during $\mathtt{T-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/3\rfloor$ distinct copies are used. Each instance of these good is allocated to a different agent. Furthermore, all the instances of the goods that are allocated during $T^{\alpha}$ along with their owners, are removed from the instance. Therefore, these goods cannot be copied more than once and their owners cannot get any other goods later on. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once and allocate the original and the duplications to $2n^{\prime}$ different agents. By Corollary 4.29, $n^{\prime}\leq\lfloor n/3\rfloor$. Therefore, in total at most $\lfloor n^{*}/3\rfloor$ distinct copies are used and the allocation is simple. ∎
C.1 Picking Sequence Fails for Allocation with Copies
@@
Consider the ordered counterpart (see Definition 4.4) of this instance and the allocation $A^{\prime}$ below, in which every good is copied once.
@@
One can generalize the picking sequence algorithm ([bouveret2016characterizing], see also Proposition 4.5), to a setting with copies in two ways. Assume we have a virtual allocation $A^{\prime}$, where each good has at most $k$ extra copies.
One way is to run $k+1$ instances of the picking sequence algorithm, where in each run each good in $M$ is allocated at most once. The first agent to pick is an agent which holds (one of the instances of) the virtual good $1$, the second holds $2$ and so on. Another way, is to allow all agents which hold good $1$ to pick first, then all agents which hold good $2$, and so on. In this interpretation agent $i$ picks the most-valued good $g$ such that $g\notin A_{i}$, and $g$ was allocated no more than $k+1$ times (so each original good is copied at most $k$ times as well). Below we show that neither of these interpretations work if we break the ties arbitrarily. In fact, the second one fails for any tie-breaking rule, see Example C.2.
The failure of the two generalizations follows from the fact that an agent is not allowed to have two instances of the same good. If valuations were such that any $q$ copies of any good $g$ would yield a value of $q\cdot v(g)$, the same proof as in [bouveret2016characterizing] would carry over to this setting. We show examples in which the two generalizations fail when at most one extra copy is made (i.e., for $k=1$). For the first generalization, the example below shows that breaking ties arbitrarily may fail.
Example C.1.